OpenStudy (anonymous):

Medals Medals! Compute the modulus and argument of each complex number.. 1. -5i 2. sqrt10(cos(5pi/7)+isin(5pi/7)

hartnn (hartnn):

modulus is 'r' argument is 'theta

hartnn (hartnn):

for 2nd, compare your expression with \(\large r(\cos \theta +i\sin \theta)\)

hartnn (hartnn):

for 1st. x+iy = 0-5i x=0, y=-5 use the formulas to calculate r and theta

OpenStudy (anonymous):

okay, so its going to look like (0, -5) right? i worked of an example i saw and my modulus became 55.9 and was stated as being "undefined"

OpenStudy (anonymous):

sorry, it was 5 and not 55.9!!

OpenStudy (anonymous):

i dont know where that came from

hartnn (hartnn):

theta was stated as undefined ?

OpenStudy (anonymous):

yeah, but im not so sure

hartnn (hartnn):

where does 0,-5 point lie ?

OpenStudy (anonymous):

on the y ais

hartnn (hartnn):

not in any quadrant :P

OpenStudy (anonymous):

yeah, my bad!

hartnn (hartnn):

 negative y axis